# Seeing the Outer Edge of the Infant Type Ia Supernova 2024epr in the Optical and Near Infrared

### v1

@@ 3.2 Color Curves
SNe 2020jgl and 2024epr are similar to the comparison SNe Ia in the color throughout their evolution. Interestingly, despite the similarity in the and colors, SNe 2020jgl and 2024epr differ from the other SNe Ia in their color curve, being much flatter ( from pre-peak to maximum light) and not showing any color evolution until 10 days past -band maximum. This implies the color difference is not solely because of reddening. One possible explanation is that, at early times, the high-velocity Ca feature is so blue-shifted that the Ca absorption trough is solely in the band (see §4.1). This difference may explain why the early-time color is much redder than the comparison SNe Ia, similar to what is seen for SN 2021aefx with its high-velocity Ca ii H&K feature (Ashall et al., 2022).
Another SN Ia with similarly low luminosity for its measured is SN 2006bt (Foley et al., 2010), which also has peculiar colors; however, the colors shown by SN 2020jgl and 2024epr ( mag) at maximum light are much redder than SN 2006bt ( mag), suggesting that SN 2024epr is unlike SN 2006bt.
@@ 4.1.1 Early-Time Comparison
On the red edge of the Si ii 6355, each SN Ia shows a potential flattening or notch feature in their earliest spectrum, which Thomas et al. (2011b) attribute to C ii 6580. The strongest feature is seen in SN 2017cbv at early times, with the C ii feature equaling the Si ii 6355 feature in depth. A clear notch is seen in SN 2011fe, with tentative C ii features seen in SNe 2020jgl and 2021aefx. SN 2024epr may have a notch similar to SN 2011fe; however, the feature’s location also corresponds to an skyline. Given the other skyline at 6800 in our day spectrum and evidence of sky subtraction issues in the subsequent two, we cannot confidently claim a C ii feature in SN 2024epr.
@@ 5 Implications for SNe Ia Models
We use these observational diagnostics to qualitatively discuss predictions made by various physical models (see sections below). While we are unable to link SN 2024epr to a single SN Ia model definitively, the high-velocity features challenge some delayed-detonation models, and we rule out He detonation models with a thick () He shell from the fits to the multiband photometry (§5.1), and the lack of He features in the NIR spectra (§5.3.1).
@@ 5.1 Model Implications from the Rising Light Curve
Some SNe Ia models may differ in their predictions for the shape of the rising light curve (e.g., Kasen 2010; Piro & Nakar 2013, 2014; Piro & Morozova 2016; Maeda et al. 2018; Polin et al. 2019; Magee & Maguire 2020; Maeda et al. 2023). Such differences may arise from intrinsic differences in the SN Ia progenitor scenario (e.g., companion interaction) or explosion mechanism (e.g., double detonation), or from an extrinsic effect like viewing angle. Observationally, there are three categories of early-time rising light curve: “single” power law rises in which the rising light curve is well-described by a single power law, “double” or “excess” rises, which are best described by a power law with an additional component (e.g., another power law or a Gaussian), and “bump” rising light curves which have non-monotonic light curves exhibiting a decline in flux before rising to the main light curve peak (Hoogendam et al., 2024).
@@
For SN 2024epr, we fit the combined rest-frame Pan-STARRS and ZTF -band light curve with a single power-law model, masking data after days. Vallely et al. (2019, 2021) find that the choice of the fitting range of a light curve can significantly change the fit results and recommend fitting only up to 40% of the SN peak flux, which, for SN 2024epr is at 10 days. Figure 13 shows the data and fit. Our model is defined as
where is a multiplicative scaling factor, is the power law index, and is the time of first light from the model. The best values from the fit are mJy , and days, or MJD . A single power-law model fits the rising light curve in this phase range well (; with an additional 2% uncertainty added to the photometry to account for potential underestimation of the photometric precision), and there is no trend in the residuals as a function of time. For those reasons, we do not perform a double power law or power law and Gaussian fit.
@@
Based on this simple power-law fit, the first optical and NIR spectra of SN 2024epr are / days after first light, respectively. This makes them remarkably early spectra for a SN Ia. Additionally, given the constraining last non-detection from ATLAS, SN Ia models that require additional flux beyond a simple single power law at early times may not match the rising light curve of SN 2024epr well.
@@
We also compare the multi-band light curves of SN 2024epr to He detonation model light curves. Figure 14 shows model He detonation absolute magnitudes from Polin et al. (2019) and our Pan-STARRS photometry. Model fits are highly sensitive to the time of first light and the distance to the SN Ia, so we perform minimization fits to determine the horizontal and vertical shifts to the model light curves to obtain the best fit for our data. Even with the phase and distance offsets (i.e., horizontal and vertical shifts to minimize the reduced value), the Polin et al. (2019) model light curves struggle to reproduce the observed light curves.
@@
Generally, the smaller 0.01 He shells match the light curves well at 10 days before peak, but the larger 0.05 He shell light curves match better at peak. No single model light curve can explain our observed data, particularly the unusual colors near peak. Given the last non-detection from ATLAS for SN 2024epr of 18.8 mag ( absolute magnitude using the redshift-inferred distance, or mag using the distance from SN 2024epr) at days before maximum light, we rule out the scenarios with larger He shells ( ) that would produce excess flux in the early-time observations. Thus, thinner He shells (e.g., Boos et al., 2021, 2024b), which do not produce excess flux in the rising light curve, may be preferential to thicker He shells to explain SN 2024epr and other SNe Ia with strong, high-velocity Ca features and smoothly rising light curves. This is not the same as suggesting that SN 2024epr definitively comes from a He-detonation, as other, non-He-detonation models may also be able to replicate the observed properties of SN 2024epr (a thorough analysis of every model is beyond the scope of this work).
@@ 5.2 Model Implications from Early-Time Optical Spectra
The presence of Si and Ca in the outermost layers of the SN 2024epr ejecta (at 38 000 km s-1) also has modeling implications. High-velocity Ca features may be from primordial Ca (Lentz et al., 2000), so such objects may be explosions of higher metallicity CO WDs. This would agree with the generally redder colors of SN 2020jgl (Galbany et al., 2025) and SN 2024epr (Section 3.2). Alternatively, the Ca may be the product of SN Ia nucleosynthesis; higher-density deflagration-to-detonation transitions in the delayed detonation models of Iwamoto et al. (1999) produce higher Ca velocities. The Iwamoto et al. (1999) models match the lack of He and the smoothly rising light curve of SN 2024epr, but it is unclear whether they can produce Ca with a velocity of 35 000 km s-1.
@@ 5.3 Model Implications from Early-Time NIR Spectra
5.3.1 Implications from the Lack of He Features
He may form spectral features in the optical. However, such potential features blend with much stronger features typical of SNe Ia, such as S ii and Si ii. In the NIR, the strongest He features, at He i 1.083 m and He i 2.058 m, suffer less severe blending.
@@ 5.3.3 Comparison to Other SNe Ia
Finally, the “notch” feature at 1.07 m seen in iPTF13ebh, SN 2022xkq, and SN 2024epr at earlier times may be Fe iii, which would suggest the presence of ionizing photons, and thus plausibly 56Ni as well, in the outermost ejecta layers. Relative to the time of the explosion, the notch appears within approximately the first two days. Fe-group elements in the outermost layers of the ejecta could arise from either enhanced 56Ni mixing or the presence of surface 56Ni. While not confirmed, understanding the origins of this notch feature may help constrain the extent of mixing or surface burning in SNe Ia.
@@ Acknowledgments
We thank Sahana Kumar, Jing Lu, Abigail Polin, Ken Shen, Eddie Baron, and Peter Hoeflich for their helpful discussions.
This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant Nos. 1842402 and 2236415. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation.
D.O.J. acknowledges support from NSF grants AST-2407632 and AST-2429450, NASA grant 80NSSC24M0023, and HST/JWST grants HST-GO-17128.028, HST-GO-16269.012, and JWST-GO-05324.031, awarded by the Space Telescope Science Institute (STScI), which is operated by the Association of Universities for Research in Astronomy, Inc., for NASA, under contract NAS5-26555. The UCSC team is supported in part by NASA grants 80NSSC23K0301 and 80NSSC24K1411; and a fellowship from the David and Lucile Packard Foundation to R.J.F. A.D. is supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant Agreement No. 101002652). A.R.W. is supported by funding through AST-2421845 and funding from the Simons Foundation for the NSF-Simons SkAI Institute, the LSST Discovery Alliance through the LINCC Incubator Program and the Catalyst Program, supported in part by Wasabi Technologies, the US DOE through the Department of Physics at the University of Illinois, Urbana-Champaign (# 13771275), and the 2024 DOE SCGSCR program and Lawrence Berkeley National Laboratory. This work was supported by National Aeronautics and Space Administration (NASA) Keck PI Data Awards, administered by the NASA Exoplanet Science Institute. This work is supported by the National Science Foundation under Cooperative Agreement PHY-2019786 (The NSF AI Institute for Artificial Intelligence and Fundamental Interactions, http://iaifi.org/).
@@
YSE-PZ was developed by the UC Santa Cruz Transients Team with support from The UCSC team is supported in part by NASA grants NNG17PX03C, 80NSSC19K1386, and 80NSSC20K0953; NSF grants AST–1518052, AST–1815935, and AST–1911206; the Gordon & Betty Moore Foundation; the Heising-Simons Foundation; a fellowship from the David and Lucile Packard Foundation to R. J. Foley; Gordon and Betty Moore Foundation postdoctoral fellowships and a NASA Einstein fellowship, as administered through the NASA Hubble Fellowship program and grant HST-HF2-51462.001, to D. O. Jones; and a National Science Foundation Graduate Research Fellowship, administered through grant No. DGE-1339067, to D. A. Coulter.

### v2

@@ 3.2 Color Curves
SNe 2020jgl and 2024epr are similar to the comparison SNe Ia in the color throughout their evolution. Interestingly, despite the similarity in the and colors, SNe 2020jgl and 2024epr differ from the other SNe Ia in their color curve, being much flatter ( mag from pre-peak to maximum light) and not showing any color evolution until 10 days past -band maximum. This implies the color difference is not solely due to reddening. One possible explanation is that, at early times, the high-velocity Ca feature is so blue-shifted that the Ca absorption trough is solely in the band (see §4.1). This difference may explain why the early-time color is much redder than that of the comparison SNe Ia, similar to what is seen for SN 2021aefx, which exhibits a high-velocity Ca ii H&K feature (Ashall et al., 2022).
Another SN Ia with similarly low luminosity for its measured is SN 2006bt (Foley et al., 2010), which also has peculiar colors; however, the colors shown by SN 2020jgl and 2024epr ( mag) at maximum light are much redder than SN 2006bt ( mag), suggesting that SN 2024epr is unlike SN 2006bt, which also had significantly slower velocities.
@@ 4.1.1 Early-Time Comparison
On the red edge of the Si ii 6355, each SN Ia shows a potential flattening or notch feature in their earliest spectrum, which Thomas et al. (2011b) attribute to C ii . The strongest feature is seen in SN 2017cbv at early times, with the C ii feature equaling the Si ii 6355 feature in depth. A clear notch is seen in SN 2011fe, with tentative C ii features seen in SNe 2020jgl and 2021aefx. SN 2024epr may have a notch similar to SN 2011fe; however, the feature’s location also corresponds to an skyline. Given the other skyline at 6800 in our day spectrum and evidence of sky subtraction issues in the subsequent two, we cannot confidently claim a C ii feature in SN 2024epr.
@@ 5 Implications for SNe Ia Models
We use these observational diagnostics to qualitatively discuss predictions made by various physical models (see sections below). While we are unable to link SN 2024epr to a single SN Ia model definitively, the high-velocity features challenge some delayed-detonation models, and we rule out He detonation models with a thick () He shell from the fits to the multiband photometry (§5.1), and the apparent lack of He features in the NIR spectra (§5.3.1).
@@ 5.1 Model Implications from the Rising Light Curve
Some SNe Ia models may differ in their predictions for the shape of the rising light curve (e.g., Kasen 2010; Piro & Nakar 2013, 2014; Piro & Morozova 2016; Maeda et al. 2018; Polin et al. 2019; Magee & Maguire 2020; Maeda et al. 2023). Such differences may arise from intrinsic differences in the SN Ia progenitor scenario (e.g., companion interaction) or explosion mechanism (e.g., double detonation), or from an extrinsic effect like viewing angle. Observationally, there are three categories of early-time rising light curve: “single” power law rises in which the rising light curve is well-described by a single power law, “double” or “excess” rises, which are best described by a power law with an additional component (e.g., another power law or a Gaussian), and “bump” rising light curves which have non-monotonic light curves exhibiting a decline in flux before rising to the main light curve peak (Hoogendam et al., 2024b, and potentially correlate with negative Hubble residuals, e.g., Ye et al., 2024).
@@
For SN 2024epr, we fit the combined rest-frame Pan-STARRS and ZTF -band light curve with a single power-law model, masking data after days. Vallely et al. (2019, 2021); Fausnaugh et al. (2021) find that the choice of the fitting range of a light curve can significantly change the fit results and recommend fitting only up to 40% of the SN peak flux, which, for SN 2024epr is at 10 days. Other studies (e.g., Fausnaugh et al., 2023; Hoogendam et al., 2024a, 2025) use a double power-law to fit transients through maximum light, negating this potential systematic. Despite having photometric coverage at this time, there is an offset in the Post light curves compared to those from ZTF and YSE, which have a negligible difference. This offset is likely due to differences in the filter functions, and it ultimately prevents us from using this more robust fitting technique. Thus, we opt to fit only at early times, with a model defined by
where is a multiplicative scaling factor, is the power law index, and is the time of first light from the model. The best values from the fit are mJy , and days, or MJD . Figure 13 shows the data and fit.
@@
A single power-law model fits the rising light curve in this phase range well (; with an additional 2% uncertainty added to the photometry to account for potential underestimation of the photometric precision). There is no trend in the residuals as a function of time. For those reasons, we do not perform a double power law or power law and Gaussian fit. Based on this simple power-law fit, the first optical and NIR spectra of SN 2024epr are / days after first light, respectively, making them remarkably early spectra for a SN Ia. Additionally, given the constraining last non-detection from ATLAS, SN Ia models that require additional flux beyond a simple single power law at early times may not accurately match the rising light curve of SN 2024epr.
@@
We also compare the multi-band light curves of SN 2024epr with those of He detonation models. Figure 14 shows model He detonation absolute magnitudes from Polin et al. (2019) and our Pan-STARRS photometry. Model fits are highly sensitive to the time of first light and the distance to the SN Ia, so we perform minimization fits to determine the horizontal and vertical shifts to the model light curves to obtain the best fit for our data. Even with the phase and distance offsets (i.e., horizontal and vertical shifts to minimize the reduced value), the Polin et al. (2019) model light curves struggle to reproduce the observed light curves.
@@
Generally, the smaller 0.01 He shells match the light curves well at 10 days before peak, but the larger 0.05 He shell light curves match better at peak. No single model light curve can explain our observed data, particularly the unusual colors near peak. Given the last non-detection from ATLAS for SN 2024epr of 18.8 mag ( absolute magnitude using the redshift-inferred distance, or mag using the distance from SN 2024epr) at days before maximum light, we rule out the scenarios with larger He shells ( ) that would produce excess flux in the early-time observations. Thus, thinner He shells (e.g., Boos et al., 2021, 2024b), which do not produce excess flux in the rising light curve, may be preferential to thicker He shells to explain SN 2024epr and other SNe Ia with strong, high-velocity Ca features and smoothly rising light curves. This is not to suggest that SN 2024epr definitively originates from a He-detonation. Non-He-detonation models may also be able to replicate the observed properties of SN 2024epr (a thorough analysis of every model is beyond the scope of this work).
@@ 5.2 Model Implications from Early-Time Optical Spectra
The presence of Si and Ca in the outermost layers of the SN 2024epr ejecta (at 38 000 km s-1) also has modeling implications. High-velocity Ca features may be from primordial Ca (Lentz et al., 2000), so such objects may be explosions of higher metallicity CO WDs. This would agree with the generally redder colors of SN 2020jgl (Galbany et al., 2025) and SN 2024epr (Section 3.2). Alternatively, the Ca may be the product of SN Ia nucleosynthesis; higher-density deflagration-to-detonation transitions in the delayed detonation models of Iwamoto et al. (1999) produce higher Ca velocities. The Iwamoto et al. (1999) models match the apparent lack of He and the smoothly rising light curve of SN 2024epr, but it is unclear whether they can produce Ca with a velocity of 35 000 km s-1.
@@ 5.3 Model Implications from Early-Time NIR Spectra
5.3.1 Implications from No Obvious He Features
He may form spectral features in the optical, but potential He features would blend with stronger features typical of SNe Ia, such as S ii and Si ii, making them difficult to detect confidently. In the NIR, however, the strongest He features, at He i 1.083 m and He i 2.058 m, suffer less severe blending.
@@ 5.3.3 Comparison to Other SNe Ia
Finally, the “notch” feature at 1.07 m seen in iPTF13ebh, SN 2022xkq, and SN 2024epr at earlier times may be Fe iii, which would suggest the presence of ionizing photons, and thus plausibly 56Ni as well, in the outermost ejecta layers. Relative to the time of the explosion, the notch appears within approximately the first two days. Fe-group elements in the outermost layers of the ejecta could arise from either enhanced 56Ni mixing or the presence of surface 56Ni (e.g., Magee et al., 2020; Magee & Maguire, 2020). While not confirmed, understanding the origins of this notch feature may help constrain the extent of mixing or surface burning in SNe Ia. One caveat to this is that a high-metallicity progenitor could also have primordial Fe that produces this feature (and may naturally also explain the redder colors), and its short lifetime could be due to an ionization change rather than a surface-heavy distribution.
@@ Acknowledgments
We thank the anonymous referee for providing helpful feedback that improved this manuscript. We thank Sahana Kumar, Jing Lu, Abigail Polin, Ken Shen, Eddie Baron, and Peter Hoeflich for their helpful discussions. This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant Nos. 1842402 and 2236415. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation. D.O.J. acknowledges support from NSF grants AST-2407632 and AST-2429450, NASA grant 80NSSC24M0023, and HST/JWST grants HST-GO-17128.028, HST-GO-16269.012, and JWST-GO-05324.031, awarded by the Space Telescope Science Institute (STScI), which is operated by the Association of Universities for Research in Astronomy, Inc., for NASA, under contract NAS5-26555. The Shappee group at the University of Hawai‘i is supported with funds from NSF (grants AST-1908952, AST-1911074, & AST-1920392) and NASA (grants HST-GO-17087, 80NSSC24K0521, 80NSSC24K0490, 80NSSC24K0508, 80NSSC23K0058, & 80NSSC23K1431). The UCSC team is supported in part by NASA grants 80NSSC23K0301 and 80NSSC24K1411; and a fellowship from the David and Lucile Packard Foundation to R.J.F. A.D. is supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant Agreement No. 101002652). A.R.W. is supported by funding through AST-2421845 and funding from the Simons Foundation for the NSF-Simons SkAI Institute, the LSST Discovery Alliance through the LINCC Incubator Program and the Catalyst Program, supported in part by Wasabi Technologies, the US DOE through the Department of Physics at the University of Illinois, Urbana-Champaign (# 13771275), and the 2024 DOE SCGSCR program and Lawrence Berkeley National Laboratory. This work is supported by the National Science Foundation under Cooperative Agreement PHY-2019786 (The NSF AI Institute for Artificial Intelligence and Fundamental Interactions, http://iaifi.org/).
This work was supported by National Aeronautics and Space Administration (NASA) Keck PI Data Awards, administered by the NASA Exoplanet Science Institute.
@@
YSE-PZ was developed by the UC Santa Cruz Transients Team, supported in part by NASA grants NNG17PX03C, 80NSSC19K1386, and 80NSSC20K0953; NSF grants AST–1518052, AST–1815935, and AST–1911206; the Gordon & Betty Moore Foundation; the Heising-Simons Foundation; a fellowship from the David and Lucile Packard Foundation to R. J. Foley; Gordon and Betty Moore Foundation postdoctoral fellowships and a NASA Einstein fellowship, as administered through the NASA Hubble Fellowship program and grant HST-HF2-51462.001, to D. O. Jones; and a National Science Foundation Graduate Research Fellowship, administered through grant No. DGE-1339067, to D. A. Coulter.